OpenStudy (anonymous):

Please help Determine whether y varies directly with x. If so, find the constant of variation k and write the equation. x y 7 11 8 13 9 15 10 17

OpenStudy (anonymous):

y does vary directly with x. as x increases by 1, y increases by 2.

OpenStudy (anonymous):

thus you know that the equation is going to be something like y = 2x + @

OpenStudy (anonymous):

plug it in anywhere, and you get y = 2x - 3

OpenStudy (anonymous):

the constant of variation is going to be 2.

OpenStudy (anonymous):

I don't understand what you mean

OpenStudy (anonymous):

look at first two lines.

OpenStudy (anonymous):

x increased from 7 to 8, right?

OpenStudy (anonymous):

and y increased from 11 to 13.

OpenStudy (anonymous):

yea I see that

OpenStudy (anonymous):

ok, lets keep that in mind and keep on going from second line to third line

OpenStudy (anonymous):

from second line to third line, x increased from 8 to 9.

OpenStudy (anonymous):

while y increased from 13 to 15.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so, from first line to second line. while x increased by 1, y increased by 2.

OpenStudy (anonymous):

from second line to third line, while x increased by 1, y increased by 2.

OpenStudy (anonymous):

this property is constantly shown in different lines.

OpenStudy (anonymous):

i see that

OpenStudy (anonymous):

"varying directly" means that the x and y changes in a constant ratio.

OpenStudy (anonymous):

in this question, x and y changes in 1:2 ratio. while x changes 1, y changes 2.

OpenStudy (anonymous):

if it's not varying directly, then the x and y will not have same ratio throughout the whole table. at one point, x will increase by 1 and y will increase by 2, when on the other point, x will increase by 1 and y might increase by 5

OpenStudy (anonymous):

okay I understand so far

OpenStudy (anonymous):

ok, first, we know that this "function" is varying directly.

OpenStudy (anonymous):

now the second part of the question asks us what the constant of variation is.

OpenStudy (anonymous):

thats what I dont understand

OpenStudy (anonymous):

ok. the constant of variation only works for functions that directly vary

OpenStudy (anonymous):

as i mentioned earlier, functions that directly vary will have a constant ratio of change between X and Y

OpenStudy (anonymous):

basically what you have to do is find how much Y is changing as X increase by 1.

OpenStudy (anonymous):

in this case, Y is increasing by 2 (constantly) as X increase by 1

OpenStudy (anonymous):

thus the constant of variation is 2.

OpenStudy (anonymous):

so do you add themtogether to get 3

OpenStudy (anonymous):

nonono. you don't add them. you have the answer already.

OpenStudy (anonymous):

constant of variation is "how much Y changes as X increases by 1"

OpenStudy (anonymous):

y=2x-3 where did the -3 come from

OpenStudy (anonymous):

ok. that's the last step.

OpenStudy (anonymous):

functions that directly change are always in one form.

OpenStudy (anonymous):

Y = K*X + B

OpenStudy (anonymous):

K is the constant of variation. in this case, it's 2

OpenStudy (anonymous):

Y and X are variables obviously, so we leave them as they are.

OpenStudy (anonymous):

so we have Y = 2X + B

OpenStudy (anonymous):

we have to figure out the B now.

OpenStudy (anonymous):

thats when we go back to the table and plug in numbers for X and Y

OpenStudy (anonymous):

im lost on this step just pug in any number

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

plug in any set of number (x,y) from the table

OpenStudy (anonymous):

as in, plug in any LINE. first line, second line, whatever.

OpenStudy (anonymous):

if you did your question correctly, then you should have the same answer for B for any line.

OpenStudy (anonymous):

lets plug in first line. it says that X = 7 and Y = 11

OpenStudy (anonymous):

we have Y = 2X + B

OpenStudy (anonymous):

we plug in 11 for Y and 7 for X

OpenStudy (anonymous):

11 = 7*2 + B

OpenStudy (anonymous):

11 = 14 + B

OpenStudy (anonymous):

thus we get B = -3

OpenStudy (anonymous):

lets try plugging in second line.

OpenStudy (anonymous):

we have X = 8 and Y = 13

OpenStudy (anonymous):

13 = 2*8 + B

OpenStudy (anonymous):

13=8*2+b b=13

OpenStudy (anonymous):

13 = 16 + B.

OpenStudy (anonymous):

uhhh.. so we have B = -3..

OpenStudy (anonymous):

not sure how you got b = 13 from that equation.

OpenStudy (anonymous):

thats what i got just didnt type it in right

OpenStudy (anonymous):

yeah i thought so.

OpenStudy (anonymous):

when you plug in the number and figure it you should get y right

OpenStudy (anonymous):

uhm. X and Y are "variables"

OpenStudy (anonymous):

meaning being, there aren't answers for them.

OpenStudy (anonymous):

i mean when you like say 13=16+b if you put-3 in the equation you end up with 13

OpenStudy (anonymous):

i think i got it thanks so much for taking tim to explain each step

OpenStudy (anonymous):

no problem.